OpenStudy (anonymous):

Thermal Dynamics - Please help me with the steps needed for the problem below, not the answer. I keep getting the wrong answer. "A construction company is making a sidewalk during winter. the average temperature during construction is 4°C. In the summer months the temperature heats up to 28°c. each sidewalk block is 1 m in length and the constant of linear expansion for the cement used is 12 x 10-6 K-1. what is the minimum amount of space between the sidewalk blocks that the constructors should use to prevent cracking during summer?"

OpenStudy (curiouser_alice):

How much will each block expand? I believe it's ∆L=α∆t, where α is the coefficient of linear expansion, and ∆t is the change in temperature. ∆t = 28-4 = 24. (This would be the same in Kelvin, since they are both positive.) Multiply by 12 to get 288 x 10 -6 m. Now fix units - =2.88 x 10 -4 m = .288 mm Next I think you would have to decide how the blocks expand - if they expand halfway on each side this should be your answer. If this answer is wrong please respond.

OpenStudy (anonymous):

WRONGGGGGGGGGGG

OpenStudy (anonymous):

it is EXPANSION=COEFFICIENT OF EXPANSION * THE FORMER LENGTH * THE CHANGE IN TEMPERATURE

OpenStudy (anonymous):

besides curiouser_alice, if it heats up by THAT MUCH and the coefficient is that LARGE, the answer is DEFINITELY WRONG